# Thinker: Learning to Think Fast and Slow

### v1

@@ 7 Future Work and Conclusion
In this paper, we introduced the Thinker task, a novel approach that decomposes the standard single-turn question-answering process into a structured four-stage task. This decomposition aims to provide more explicit reward signals for training distinct agent capabilities: intuition via Fast Thinking, evaluation via Verification, refinement via Slow Thinking, and integration via Summarization. Our experiments demonstrate significant performance improvements across a majority of mathematical reasoning benchmarks. Furthermore, our analysis indicates that agents trained with the Thinker task exhibit a reduction in overt reflection patterns during their reasoning process, suggesting more direct problem-solving.
Beyond the specific application to question-answering, this work underscores the potential of environment augmentation in RL. In RL, while significant attention is often devoted to algorithm development, the environment itself is typically treated as a given problem specification. However, as demonstrated by this paper and prior research such as the Thinker MDP, there is considerable untapped potential in designing environments that offer richer inputs, more structured interactions, or more nuanced reward signals. Future research could explore alternative methods for enriching RL environments, perhaps by developing more dynamic tasks that adapt to an agent’s learning state or that explicitly target the development of a wider array of cognitive capabilities. Such advancements in environment design could unlock new levels of performance and generalization in RL agents.
Acknowledgement
Stephen Chung completed this work at the University of Cambridge, while Wenyu Du completed this work at the University of Hong Kong.
@@ Appendix A Experimental Details
This section describes the details of the experiment in Section 5. Our implementation, adapted from the Open-Reasoner-Zero codebase [3], is publicly available at https://github.com/DualityRL/thinker-task.
@@ A.1 Hyperparameters
[…]
@@ B.1 Evaluation Results
The detailed evaluation results of the ablated run in which the Summarization step is removed can be found in Table 3, labeled as SkipSum. Table 4 presents the standard error corresponding to the results reported in Table 1.
@@
Figure 7: Detailed evaluation results of Q1.5B fine-tuned using QA task or Thinker task on individual mathematical reasoning benchmarks.
@@
Table 3: Mathematical reasoning performance of the Thinker agent trained without the Summarization stage. Average (Avg.) scores are presented. All scores are Pass@1 accuracy (%) averaged over 16 samples. Standard errors are provided in Table 4.
@@
SkipSum 60.30 5.00 1.25 20.27 24.17 30.00 19.85 38.24 24.88
SkipSum-Fast 64.30 9.17 4.17 18.62 29.11 37.50 20.82 39.42 27.89
Table 4: Standard error of performance on mathematical benchmarks in Table 1. All scores are in %.
Method MATH AIME AIME GPQA Olympiad AMC Minerva College Avg.
500 2024 2025 Diamond bench 23 Math Math
Qwen2.5-1.5B (Q1.5B)
Pretrained 0.36 0.00 0.00 0.40 0.15 0.91 0.19 0.09 0.14
Baseline 0.33 0.61 0.86 0.47 0.34 1.20 0.50 0.11 0.24
Thinker 0.19 0.67 0.57 0.39 0.09 1.32 0.13 0.08 0.17
Thinker-Fast 0.19 0.52 0.57 0.46 0.27 0.85 0.43 0.11 0.10
SkipSum 0.23 0.61 0.42 0.91 0.21 1.25 0.48 0.05 0.20
SkipSum-Fast 0.39 0.57 0.71 0.51 0.22 1.37 0.22 0.09 0.12
DeepSeek-R1-Distill-Qwen-1.5B (R1.5B)
Pretrained 0.33 1.34 0.96 0.58 0.28 1.20 0.42 0.08 0.24
Baseline 0.24 1.05 0.52 0.56 0.34 0.99 0.30 0.08 0.22
Thinker 0.25 1.35 1.09 0.63 0.31 0.96 0.38 0.09 0.16
Thinker-Fast 0.25 1.10 0.91 0.45 0.21 1.46 0.24 0.07 0.24
@@ B.2 Ablation Study on Fast Thinking Mode
To understand the importance of the Fast Thinking mode in the overall Thinker task, we experiment by using a less-trained agent to generate the Fast Thinking response, while still using the fully trained agent to generate responses for the remaining stages. This allows us to measure the impact of Fast Thinking quality on overall performance.
@@ Appendix C Case Study
Fast Thinking. The agent adopts a quick, System-1-like heuristic despite expressed uncertainty (“the side length of the larger hexagon is $a+2\cdot 2=a+4$? Not sure.”) and proceeds with this flawed assumption ($S=a+4$). This leads to an incorrect perimeter, $126\sqrt{3}-$ 12, which is physically implausible as it results in a negative value.
Verification. The agent directly confronts this flawed assumption. It explicitly questions the initial logic (“The side length of this larger hexagon might be $s+4$? Or is it $s+2$? Wait, reconsider. <…> the relationship isn’t straightforward.”). This critical re-evaluation leads to the correct geometric insight, yielding the relationship $S=s+\frac{\sqrt{3}}{4}$. Based on this corrected understanding, the agent identifies the error in its initial reasoning and conclusion, stating, “The previously given answer was $126\sqrt{3}-$ 12. But that would not match. <…> Thus our initial approach is wrong.”
@@
Box C.1: Example Response from Thinker Agent: Hexagon Path Problem Problem: A path $2\mathrm{~{}m}$ wide and with an area of $36\mathrm{~{}m}^{2}$ has been built around a swimming pool in the shape of a regular hexagon. What is the perimeter of the pool? User: <Fast Thinking Prompt: Answer concisely…> Assistant: <Initial Reasoning Trace showing uncertainty about side length> 3. <…> the side length of the larger hexagon is $a+2\cdot 2=a+4$? Not sure. <Calculations based on the assumption $S=a+4$> 13. So $28a=8\sqrt{3}-16\Rightarrow a=\sqrt{3}-$ 2. 14. Perimeter of the hexagon is $126a=6(\sqrt{3}-2)=6\sqrt{3}-$ 12. $\boxed{6\sqrt{3}-12}$ User: <Verification Prompt: Is your answer above correct?> Assistant: <Reasoning Trace in Verification, questioning the $S=s+4$ assumption ($a$ relabeled as $s$)> The side length of this larger hexagon might be $s+4$? Or is it $s+2$? Wait, reconsider. <…> the relationship isn’t straightforward. <Derivation of the correct relationship $S=s+\frac{4}{\sqrt{3}}$ based on apothems> <Calculations using the corrected $S$, leading to Perimeter = $18-4\sqrt{3}$> The previously given answer was $126\sqrt{3}-$ 12. But that would not match. So our approach seems incorrect. <…> Thus our initial approach is wrong. Therefore, the original answer is incorrect. <…> $\boxed{No}$ User: <Slow Thinking Prompt: Your initial answer is incorrect. Now, think about errors…> Assistant: <think> <revisiting the core geometric error from Fast Thinking> When you add a path around the hexagon, the increase in side length can be found by considering the distance from the center. <…> The path is 2m wide, which I think refers to the distance from the pool to the path, that would mean the difference in the apothems is 2m. So: $\frac{S\sqrt{3}}{2}-\frac{s\sqrt{3}}{2}=2$ <Re-derivation of $S=s+\frac{4}{\sqrt{3}}$, consistent with Verification’s insight> $S=s+\frac{4}{\sqrt{3}}$ <Systematic algebraic solution using the corrected $S$, leading to $s=3-\frac{2\sqrt{3}}{3}$> <Calculation of perimeter = $18-4\sqrt{3}$, followed by numerical verification> Hence, the perimeter is $6s\approx 18-4\sqrt{3}$ meters. <…> Yes, that seems consistent. Final Answer. The perimeter of the pool is $\boxed{18-4\sqrt{3}}$ meters. </think>
@@
2. The area of the path is the difference between the area of the larger hexagon and the smaller one. The area of a regular hexagon is $(3\sqrt{3}/2)\cdot(\text{side length})^{2}$.
@@
Let $s$ be the side length of the original hexagonal pool. The path width is $2\mathrm{m}$. The side length of the larger hexagon (pool + path) is $S=s+\frac{4}{\sqrt{3}}$. The area of the path is $36\mathrm{~{}m}^{2}$, which is the difference between the areas of the larger and smaller hexagons. The area of a regular hexagon is $\frac{3\sqrt{3}}{2}\times(\text{side length})^{2}$. Therefore: $\frac{3\sqrt{3}}{2}\left(\left(s+\frac{4}{\sqrt{3}}\right)^{2}-s^{2}\right)=36$ Simplifying the term inside the brackets: $\left(s+\frac{4}{\sqrt{3}}\right)^{2}-s^{2}=s^{2}+\frac{8s}{\sqrt{3}}+\frac{16}{3}-s^{2}=\frac{8s}{\sqrt{3}}+\frac{16}{3}$ Substituting this back into the area equation: $\frac{3\sqrt{3}}{2}\left(\frac{8s}{\sqrt{3}}+\frac{16}{3}\right)=36$ Multiplying through by $\frac{3\sqrt{3}}{2}$: $\frac{3\sqrt{3}}{2}\cdot\frac{8s}{\sqrt{3}}+\frac{3\sqrt{3}}{2}\cdot\frac{16}{3}=36$ $3612s+8\sqrt{3}=$ 36 Solving for $s$: $12s=36-8\sqrt{3}$ $s=\frac{36-8\sqrt{3}}{12}=3-\frac{2\sqrt{3}}{3}$ The perimeter of the pool is $6s$: $6\left(3-\frac{2\sqrt{3}}{3}\right)=18-4\sqrt{3}$ Thus, the perimeter of the pool is $\boxed{18-4\sqrt{3}}$ meters.
Box C.3: Example Response from Thinker Agent: Sequence Product Problem Problem: Define the sequence $\left\{a_{n}\right\}:a_{1}=1,a_{2}=2,a_{n+2}=a_{n}+a_{n+1},n\in\mathbf{N}_{+}$, then find $\left[\frac{a_{2}}{a_{1}}\right]\cdot\left\{\frac{a_{3}}{a_{2}}\right\}\cdot\ldots\cdot\left\{\frac{a_{99}}{a_{98}}\right\}\cdot\left[\frac{a_{98}}{a_{2}}\right]$. Options: A. 2, B. 0, C. 1, D. $\frac{1}{2}$ User: <Fast Thinking Prompt: Answer concisely…> Assistant: <Initial computation of sequence terms and first few product terms: $[a_{2}/a_{1}]=2$, $\{a_{3}/a_{2}\}=1/2$> 7. Notice that the product of fractional parts potentially telescopes or simplifies. However, it’s complex. 8. Given the options A to D, and considering the pattern, likely answer is B.0 or C.1, but need to think carefully. Since initial term is 2 and product involves fractions potentially leading to 1/2. $\boxed{D}$ User: <Verification Prompt: Is your answer above correct?> Assistant: <Re-evaluation of sequence and product structure, calculating $[a_{2}/a_{1}]=2$ and $[a_{98}/a_{2}]\approx 17$> Wait, count the number of fractional parts: <…> 96 terms. <…> Since each term $\{a_{k}/a_{k-1}\}$ is less than 1, and multiple such terms multiplied, so perhaps the total product tends to zero. Thus, the product is very small, so the overall expression would be close to zero. <…> Therefore, the original answer was incorrect, so the correct answer should be $\boxed{No}$, as the initial answer was incorrect. <…> now I think the final answer might be B.0. User: <Slow Thinking Prompt: Your initial answer is incorrect. Now, think about errors…> Assistant: <think> <Recalculates sequence terms, $[a_{2}/a_{1}]=2$. Identifies $a_{n}=F_{n+1}$ and that $F_{99}$ is even, refining $[a_{98}/a_{2}]$ to $F_{99}/2$> Given that each ratio $a_{k}/a_{k-1}$ approaches the golden ratio $\phi\approx 1.618$. Thus, each $\{a_{k}/a_{k-1}\}$ approaches $\phi-1=\psi\approx 0.618$. Thus, if each term $\{a_{k}/a_{k-1}\}$ is about $\psi\approx 0.618$. Thus, product of 96 terms [error in count, should be 97]: $\psi^{96}$. Therefore, product=$2\cdot(\psi)^{96}\cdot\text{floor}(F_{99}/2)$. <…> Alternatively, since the product seems to be approaching zero due to the product of many terms each less than 1. But given the options, perhaps it’s 0. Therefore, the answer is B.0. $\boxed{B}$ </think>
@@
Appendix D Preliminary Experiments on Larger Models
To understand the scalability of the Thinker task, we fine-tune DeepSeek-R1-Distill-Qwen-7B (R7B) [1], a larger version of R1.5B that also undergoes distillation using reasoning data from DeepSeek-R1. We use the same set of hyperparameters and training data as in the main experiments. The baseline corresponds to fine-tuning R7B on the standard QA task.
The average evaluation accuracy across the seven benchmarks (excluding CollegeMath due to the large number of questions) during training is shown in Figure 10. Due to limited computational resources, we have not yet trained the Thinker agent to convergence, but we already observe its performance surpassing that of the baseline, which has plateaued. The detailed performance of the best checkpoints from both runs can be found in Table 5.
Similar to the results observed in smaller models, we found that agents trained under the Thinker task consistently perform better than those trained on the baseline QA task across all benchmarks. The overall performance improved from 54.41% to 58.91%, representing a relative improvement of 8.27%. This suggests that larger models also benefit from the Thinker task. Additionally, we note that Thinker-Fast performance improves from 37.55% in R1.5B to 45.45% in R7B, demonstrating that the Fast Thinking mode scales well with model size.

### v2

@@ 6 Future Work and Conclusion
In this paper, we introduced the Thinker task, a novel approach that decomposes the standard single-turn question-answering process into a structured four-stage task. This decomposition aims to provide more explicit reward signals for training distinct agent capabilities: intuition via Fast Thinking, evaluation via Verification, refinement via Slow Thinking, and integration via Summarization.
Beyond the application to question-answering, this work underscores the potential of environment augmentation in RL. In RL, while significant attention is devoted to algorithm development, the environment itself is typically treated as a given problem specification. However, as demonstrated by this paper and prior research such as the Thinker MDP, there is considerable untapped potential in designing environments that offer richer inputs, more structured interactions, or more nuanced reward signals. Future research could explore alternative methods for enriching RL environments, perhaps by developing more dynamic tasks that adapt to an agent’s learning state or that explicitly target the development of a wider array of cognitive capabilities. Such advancements in environment design could unlock new levels of performance and emergent abilities in RL agents.
@@ Appendix A Experimental Details
This section describes the details of the experiment in Section 5. Our implementation, adapted from the Open-Reasoner-Zero codebase [3], is publicly available at https://github.com/stephen-chung-mh/thinker-task, which also includes the trained models.
@@ A.1 Hyperparameters
We tune the coefficient $c$ in $R_{\text{summary}}$ by searching over $\{1\text{e-}4,1\text{e-}3,1\text{e-}2\}$. Other Thinker-task-specific hyperparameters are selected using heuristics. We use a lower sampling temperature during summarization, as we observe that higher temperatures tend to produce less concise and consistent summaries.
@@ B.1 Evaluation Results
The detailed evaluation results of the ablated run in which the Summarization step is removed can be found in Table 4, labeled as SkipSum.
Table 5 and Table 6 present the standard error corresponding to the results reported in Table 1. The standard errors here are computed using Equation (1) from [30].
@@
Figure 8: Detailed evaluation results of Q1.5B fine-tuned using QA task or Thinker task on individual mathematical reasoning benchmarks. The shaded region represents the standard deviation across three independent seeds
@@
Table 4: Mathematical reasoning performance of the Thinker agent trained without the Summarization stage. Average (Avg.) scores are presented. All scores are Pass@1 accuracy (%) averaged over 16 samples. All runs used a 6000-token Verification step.
@@
Thinker 64.25 6.25 2.50 23.74 28.11 40.62 19.03 38.33 27.85
Thinker-Fast 61.60 6.25 2.50 26.39 24.78 35.94 18.66 37.85 26.75
SkipSum 64.30 9.17 4.17 18.62 29.11 37.50 20.82 39.42 27.89
SkipSum-Fast 60.30 5.00 1.25 20.27 24.17 30.00 19.85 38.24 24.88
Table 5: Standard error analysis for the Q1.5BB model. All scores are in %. The values are presented as score (standard error).
Benchmark # Qs Thinker Thinker-Fast Baseline
Seed 1 Seed 2 Seed 3 Seed 1 Seed 2 Seed 3 Seed 1 Seed 2 Seed 3
MATH 500 500 63.25 (1.83) 64.72 (1.79) 65.38 (1.80) 59.42 (1.83) 59.67 (1.85) 60.38 (1.84) 57.98 (1.79) 60.72 (1.80) 60.75 (1.79)
AIME 2024 30 5.62 (2.51) 5.83 (3.20) 7.29 (3.34) 5.00 (2.83) 4.17 (2.69) 4.58 (3.16) 3.33 (1.58) 4.38 (2.61) 4.58 (2.49)
AIME 2025 30 3.54 (1.84) 1.88 (1.13) 1.25 (0.70) 1.46 (1.26) 1.04 (0.85) 1.25 (0.63) 3.33 (1.79) 1.46 (1.07) 2.50 (1.36)
GPQA Diamond 198 19.00 (1.69) 17.99 (1.67) 20.64 (1.78) 23.11 (1.77) 19.85 (1.84) 20.86 (1.81) 21.46 (1.80) 19.35 (1.61) 20.74 (1.74)
Olympiad Bench 675 28.02 (1.44) 27.94 (1.45) 28.67 (1.46) 24.82 (1.35) 24.13 (1.37) 24.60 (1.35) 24.54 (1.32) 26.46 (1.39) 27.16 (1.38)
AMC23 40 38.91 (6.11) 41.41 (6.38) 36.88 (6.18) 36.09 (6.43) 34.84 (5.74) 32.66 (5.78) 34.38 (5.62) 36.41 (5.67) 35.31 (5.46)
Minerva Math 272 19.90 (1.92) 21.83 (2.02) 19.42 (1.90) 16.98 (1.75) 20.06 (1.90) 16.50 (1.74) 17.78 (1.85) 19.88 (1.88) 20.08 (1.94)
College Math 2818 38.73 (0.85) 39.15 (0.86) 38.58 (0.85) 37.49 (0.85) 37.84 (0.84) 37.39 (0.84) 35.85 (0.81) 38.26 (0.84) 38.15 (0.85)
Table 6: Standard error analysis of R1.5B models on mathematical benchmarks. All scores are in %. The values are presented as score (standard error).
Benchmark # Questions Thinker Thinker-Fast Baseline
MATH 500 500 88.51 (1.22) 81.35 (1.50) 86.24 (1.25)
AIME 2024 30 38.96 (7.21) 18.33 (5.23) 35.42 (7.09)
AIME 2025 30 26.67 (7.52) 14.58 (5.37) 23.75 (6.74)
GPQA Diamond 198 37.41 (2.24) 28.85 (2.27) 25.69 (2.00)
Olympiad Bench 675 55.49 (1.67) 45.68 (1.66) 49.22 (1.66)
AMC23 40 83.59 (4.70) 66.41 (5.94) 72.81 (5.18)
Minerva Math 272 34.77 (2.57) 31.39 (2.43) 32.08 (2.47)
College Math 2781 42.46 (0.90) 41.74 (0.89) 42.02 (0.89)
@@ B.2 Ablation Study on Fast Thinking Mode
To understand the importance of the Fast Thinking mode in the overall Thinker task, we experiment by using a less-trained agent to generate the Fast Thinking response, while still using the fully trained agent to generate responses for the remaining stages. [^2] The section here used a 6000-token Verification step. The length was later reduced to 2000 tokens for the main experiments, as this did not significantly affect performance. This allows us to measure the impact of Fast Thinking quality on overall performance.
@@ Appendix C Case Study
Fast Thinking. The agent adopts a quick, System-1-like heuristic despite expressed uncertainty (“the side length of the larger hexagon is $a+2\cdot 2=a+4$? Not sure.”) and proceeds with this flawed assumption ($S=a+4$). This leads to an incorrect perimeter, $6\sqrt{3}-12$, which is physically implausible as it results in a negative value.
Verification. The agent directly confronts this flawed assumption. It explicitly questions the initial logic (“The side length of this larger hexagon might be $s+4$? Or is it $s+2$? Wait, reconsider. <…> the relationship isn’t straightforward.”). This critical re-evaluation leads to the correct geometric insight, yielding the relationship $S=s+\frac{\sqrt{3}}{4}$. Based on this corrected understanding, the agent identifies the error in its initial reasoning and conclusion, stating, “The previously given answer was $6\sqrt{3}-12$. But that would not match. <…> Thus our initial approach is wrong.”
@@
Box C.1: Example Response from Thinker Agent: Hexagon Path Problem Problem: A path $2\mathrm{~m}$ wide and with an area of $36\mathrm{~m}^{2}$ has been built around a swimming pool in the shape of a regular hexagon. What is the perimeter of the pool? User: <Fast Thinking Prompt: Answer concisely…> Assistant: <Initial Reasoning Trace showing uncertainty about side length> 3. <…> the side length of the larger hexagon is $a+2\cdot 2=a+4$? Not sure. <Calculations based on the assumption S=a+4S=a+4> 13. So $8a=8\sqrt{3}-16\Rightarrow a=\sqrt{3}-2$. 14. Perimeter of the hexagon is $6a=6(\sqrt{3}-2)=6\sqrt{3}-12$. $\boxed{6\sqrt{3}-12}$ User: <Verification Prompt: Is your answer above correct?> Assistant: <Reasoning Trace in Verification, questioning the $S=s+4$ assumption ($a$ relabeled as ss)> The side length of this larger hexagon might be $s+4$? Or is it $s+2$? Wait, reconsider. <…> the relationship isn’t straightforward. <Derivation of the correct relationship $S=s+\frac{4}{\sqrt{3}}$ based on apothems> <Calculations using the corrected $S$, leading to Perimeter = $18-4\sqrt{3}$> The previously given answer was $6\sqrt{3}-12$. But that would not match. So our approach seems incorrect. <…> Thus our initial approach is wrong. Therefore, the original answer is incorrect. <…> $\boxed{No}$ User: <Slow Thinking Prompt: Your initial answer is incorrect. Now, think about errors…> Assistant: <think> <revisiting the core geometric error from Fast Thinking> When you add a path around the hexagon, the increase in side length can be found by considering the distance from the center. <…> The path is 2m wide, which I think refers to the distance from the pool to the path, that would mean the difference in the apothems is 2m. So: $\frac{S\sqrt{3}}{2}-\frac{s\sqrt{3}}{2}=2$ <Re-derivation of $S=s+\frac{4}{\sqrt{3}}$, consistent with Verification’s insight> $S=s+\frac{4}{\sqrt{3}}$ <Systematic algebraic solution using the corrected $S$, leading to $s=3-\frac{2\sqrt{3}}{3}$> <Calculation of perimeter = $18-4\sqrt{3}$, followed by numerical verification> Hence, the perimeter is $6s\approx 18-4\sqrt{3}$ meters. <…> Yes, that seems consistent. Final Answer. The perimeter of the pool is $\boxed{18-4\sqrt{3}}$ meters. </think>
@@
[…]
@@
Let $s$ be the side length of the original hexagonal pool. The path width is $2\mathrm{m}$. The side length of the larger hexagon (pool + path) is $S=s+\frac{4}{\sqrt{3}}$. The area of the path is $36\mathrm{~m}^{2}$, which is the difference between the areas of the larger and smaller hexagons. The area of a regular hexagon is $\frac{3\sqrt{3}}{2}\times(\text{side length})^{2}$. Therefore: $\frac{3\sqrt{3}}{2}\left(\left(s+\frac{4}{\sqrt{3}}\right)^{2}-s^{2}\right)=36$ Simplifying the term inside the brackets: $\left(s+\frac{4}{\sqrt{3}}\right)^{2}-s^{2}=s^{2}+\frac{8s}{\sqrt{3}}+\frac{16}{3}-s^{2}=\frac{8s}{\sqrt{3}}+\frac{16}{3}$ Substituting this back into the area equation: $\frac{3\sqrt{3}}{2}\left(\frac{8s}{\sqrt{3}}+\frac{16}{3}\right)=36$ Multiplying through by $\frac{3\sqrt{3}}{2}$: $\frac{3\sqrt{3}}{2}\cdot\frac{8s}{\sqrt{3}}+\frac{3\sqrt{3}}{2}\cdot\frac{16}{3}=36$ $12s+8\sqrt{3}=36$ Solving for $s$: $12s=36-8\sqrt{3}$ $s=\frac{36-8\sqrt{3}}{12}=3-\frac{2\sqrt{3}}{3}$ The perimeter of the pool is $6s$: $6\left(3-\frac{2\sqrt{3}}{3}\right)=18-4\sqrt{3}$ Thus, the perimeter of the pool is $\boxed{18-4\sqrt{3}}$ meters.
[…]
@@
Appendix D Additional Experiments on Larger Models
To assess the scalability of the Thinker task, we fine-tuned DeepSeek-R1-Distill-Qwen-7B (R7B) [1], a larger version of the R1.5B model that has also undergone distillation using reasoning data from DeepSeek-R1. The same set of hyperparameters and training data from the main experiments were utilized, with the exception of an increased token budget for the verification stage—from 2,000 to 6,000 tokens—to accommodate the more verbose reasoning of larger models. The baseline for this experiment corresponds to fine-tuning the R7B model on the standard QA task.
The average evaluation accuracy across the seven benchmarks (excluding CollegeMath) during training is shown in Figure 11. The detailed performance of the best checkpoints from both runs can be found in Table 7. Similar to the results observed in smaller models, we found that agents trained under the Thinker task consistently perform better than those trained on the baseline QA task across all benchmarks. The overall performance improved from 54.41% to 59.09%, representing a relative improvement of 8.6%. This suggests that larger models also benefit from the Thinker task. Additionally, we note that Thinker-Fast performance improves from 41.05% in R1.5B to 47.19% in R7B, demonstrating that the Fast Thinking mode scales well with model size.